SpaceRef

SpaceRef


NASA Space Station On-Orbit Status 4 Jun 2004

Status Report From: NASA HQ
Posted: Friday, June 4, 2004

image All ISS systems continue to function nominally except those noted previously or below.

CDR Gennady Padalka started preparations of Orlan equipment for the upcoming EVA (still envisioned for 6/15 or 6/16, but possibly slipping to 6/28 to support a more favorable crew sleep cycle).   [Gennady's work today dealt with gathering and configuring Orlan systems, setting up the ZU-S battery charger, and initiating charging on the 825M3 batteries (28V) for the Orlan backpacks.  The BITS2-12 telemetry system, VD-SU control mode, Elektron unit and SKV-1 air conditioner had to be temporarily deactivated for setting up the cable connections for the charging.  Correction of yesterday's status: The BNP repress bottles installed by Padalka in DC1 and SM contain air, not oxygen.]

FE/SO Michael Fincke meanwhile reviewed new EVA procedures and started gathering necessary tools and equipment.  He also continued the NiMH battery discharge/recharge maintenance activities, initiating another servicing cycle on four EHIP (EMU helmet interchangeable portable) batteries and three PGT (pistol grip tool) batteries.  Termination of these activities will depend on the actual EVA date finally selected.   [The EHIP batteries are numbered 1009, 1016, 1018 & 1020; the PGT batteries are 1005, 1018 & 1015.  It is planned to attach U.S.EMU helmet lights to the Orlans (by "wire ties") for facilitating the nighttime work.  Procedures uplinked for crew review include a list of hand signals to be used in the event of comm loss between the Orlans, use of the GStM "Strela" cargo boom that will be extended from its base at the DC1 and attached at a GStM grapple fixture on the FGB, etc.]

Mike Fincke retrieved the spare RPCM (remote power controller module), scheduled to replace the failed RPCM on the S0 truss, from its stowage behind the one of the Lab ZSRs (zero-G storage racks).  This required him to unbolt and rotate the ZSR out of the way.   [To prevent structural interferences during the rotation, a smoke detector and ARIS (active rack isolation system) microgravity rack barrier had to be removed first.]

Continuing the current round of the monthly preventive maintenance of RS air ventilation systems, the CDR today cleaned the Group B ventilation fans in the SM, including a checkout of the ventilators and a thorough cleaning of fan screens and flexible air ducts.  Before the cleaning, he took photographs of specific fan grids for the ground to inspect their condition.

The FE completed the scheduled full calibration of CSA-CP (compound specific analyzer-combustion products) #1010, using MCA (major constituents analyzer) measurements as calibration reference.  To ensure accurate MCA readings, the ground commanded a full calibration of the MCA after it had been operating for ~6 hrs, supported by the crew by opening and closing the manual oxygen valve (HVO2) before and after the calibration.  [#1010 is prime for combustion products and will not be used for oxygen measurement; the latter will be done with a second CSA-CP, #1003, whose O2 sensor will be calibrated separately.  For O2 measurements, MCA will be activated for comparison.]

Afterwards, Mike resupplied the CSA-CP stowage kit with new filters delivered by 14P and any unused lithium batteries from the 12P CDMK (CO2 monitoring kit).   [The two new CSA-CPs delivered on 14P (#1015, #1016) will need to be deployed and left in the open atmosphere for several weeks of decontamination before they are operational.]

Fincke performed the daily leak check on the Lab window pressure test equipment.

The CDR completed the daily routine maintenance of the SM's SOZh life support system, and prepared the daily "delta" file for updating the IMS (inventory management system) database(s).  The Science Officer conducted the regular routine status checkup of autonomously operating Increment 9 payloads.

Gennady held his weekly IMS tagup with ground specialists, discussing open issues concerning identification of equipment and storage locations for updating the IMS database, via S-band.

The Elektron O2 generator continues to operate nominally (18A mode).  Fincke performed another repressurization of its BZh-7 liquid unit, which is known to have a tiny leak inside the sealed container.  As a regular procedure, whenever the inside pressure reaches 1.0 atm the unit is repressed with nitrogen (to prevent outside air getting in).

At about 2:15pm EDT, the crew's schedule calls for their weekly teleconference with the ISS Flight Director at MCC-H.

Today's CEO (Crew Earth Observations) photo targets, limited in XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in "ram"), were Internal waves, Aegean Sea (looking a touch right towards the near-noon sunglint point and shooting any internal waves that may pass within the glint disc), Algiers, Algeria (nadir pass over this port city), Internal waves, Bahamas (looking a touch right near the glint point for any packets of internal waves generated between the islands), Los Angeles, CA (looking left for the LA basin), and  Las Vegas, NV (looking a touch left for the desert city).

CEO images can be viewed at the websites.

See also the website "Space Station Challenge" at

U.S. and Russian Segment Status

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is now functioning again).  SFOG slot #2 fan suspect (not usable).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (solar-tracking, "sun slicer", i.e., drag reduction-biased by 47 deg).
  • SM batteries:  Battery #3 is on line in "Cycle" mode.  All other batteries (7) are on line in "Partial Charge" mode.  
  • FGB batteries:  Battery #5 is off line; all other batteries (5) are on line in "Partial Charge" mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3436 kg (7575 lb) as of 5/20/04;  [SM(552) + FGB(2884) + Progress M-1(0)].  (Capability: SM -- 860 kg; FGB -- 6120 kg).

Attitude Control Systems:

  • 2 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2's RPC-17 failed 4/21/04).
  • State vector source -- U.S. SIGI-1 (GPS)
  • Attitude source -- U.S. SIGI-1 (GPS)
  • Angular rate source -- RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = "sun-fixed" [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 5:32am EDT [= epoch]):

  • Mean altitude -- 362.6 km
  • Apogee -- 366.6 km
  • Perigee -- 358.7 km
  • Period -- 91.8 min.
  • Inclination (to Equator) --  51.6336 deg
  • Eccentricity -- 0.0005835
  • Solar Beta Angle --  -13.6 deg
  • Orbits per 24-hr. day -- 15.69
  • Mean altitude loss in last 24 hours -- 65 m
  • Revolutions since FGB/Zarya launch (Nov. '98)  -- 31630

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see  http://www.hq.nasa.gov/osf/station/viewing/issvis.html

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.